OpenStudy (anonymous):

Find the product of (4x + 3y)(4x − 3y).

OpenStudy (anonymous):

it is 16x^2 - 9y^2

OpenStudy (anonymous):

because we know (a+b)(a-b) = (a^2 -b^2) in this case a=4x, b=3y :)

OpenStudy (anonymous):

thanks =)